OpenStudy (anonymous):

How does 1 1/3 = 1 2/6? I'm starting to learn about fractions more because it's something I missed a lot of in grad school and now I'm trying to catch up. I learned that the dominator is 6 since it's decided up into 6 equal parts but how can 1 1/3 come out of that?

OpenStudy (anonymous):

Because when you simplify you are able to divide 2 into both the denominator and the numerator.

OpenStudy (amistre64):

you know that 1 = 1 so that part is not a problem right? :)

OpenStudy (amistre64):

if you have 2 pies; and cut one into 3 equal pieces and the other into 6 equal pieces; when you compare the 2 pies you see that: 2 pieces out of the 6 slices equals the same amount as the 1 slice out of the 3 pieces...

OpenStudy (amistre64):

1 out of 3 = 2 out of 6 1/3 = 2/6

OpenStudy (amistre64):

we can simplify this by simply saying that: 1/3 of a pis is 1/3 of a pis no matter how many pieces the pie is cut into

OpenStudy (amistre64):

gotta read thru the typos lol

OpenStudy (amistre64):

4 pieces out of 12 slices is still 1/3 of the pie 4/12 = 1/3

OpenStudy (anonymous):

lol, that's okay, it's just a bit hard to grasp for me, I don't know why that is exactly but I'm thinking over what you're saying, I get a lot of it but somehow I still have a hard time understanding it

OpenStudy (amistre64):

6 pieaces out of 18 slices is still 1/3 of the pie 6/18 = 1/3

OpenStudy (amistre64):

we can try it with money :) one quarter = 1 fourth of a dollar 1/4 = 1/4 25 pennies = 1/4 of a dollar 25/100 = 1/4

OpenStudy (anonymous):

Okay, I'm with you lol

OpenStudy (amistre64):

:)

OpenStudy (radar):

When it comes to money, they understand lol.

OpenStudy (amistre64):

:)